# Image Intrinsic Scale Assessment: Bridging the Gap Between Quality and Resolution

### v2

@@ Abstract
Image Quality Assessment (IQA) measures and predicts perceived image quality by human observers. Although recent studies have highlighted the critical influence that variations in the scale of an image have on its perceived quality, this relationship has not been systematically quantified. To bridge this gap, we introduce the Image Intrinsic Scale (IIS), defined as the largest scale where an image exhibits its highest perceived quality. We also present the Image Intrinsic Scale Assessment (IISA) task, which involves subjectively measuring and predicting the IIS based on human judgments. We develop a subjective annotation methodology and create the IISA-DB dataset, comprising 785 image-IIS pairs annotated by experts in a rigorously controlled crowdsourcing study with verified reliability. Furthermore, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a strategy that leverages how the IIS of an image varies with downscaling to generate weak labels. Experiments show that applying WIISA during the training of several IQA methods adapted for IISA consistently improves the performance compared to using only ground-truth labels. We will release the code, dataset, and pre-trained models upon acceptance.
* Equal contribution.$\dagger$ Work done during an internship at Sony AI.
@@ 1 Introduction
Most IQA datasets annotate quality ratings – in the form of Mean Opinion Scores (MOS) – for images at a fixed resolution [14, 15, 9]. To the best of our knowledge, KonX [33] was the first dataset that provided quality annotations for each image rescaled across multiple resolutions. The authors showed that perceived quality varies with image resolution (i.e., size in pixels). When an image is downscaled, its resolution decreases. Since images are typically displayed at a consistent pixel density on any given display medium, downscaled images also appear physically smaller. The perceived quality of smaller images often increases because degradation becomes less noticeable, e.g., noise grain becomes smaller or fades. Conversely, relevant high-frequency details may get lost when the resolution is too low, leading to relatively lower perceived quality. The top section of Fig. 1 shows an example of the former phenomenon. The image at the original size ($scale=1$) appears blurry and degraded. However, when downscaled ($scale=0.25$), the degradations are not noticeable, improving quality, while relevant details, such as the bird’s feathers and barbs, are crisp and clear. At smaller scales ($scale=0.1$), the individual barbs cannot be resolved anymore, and detail is lost.
@@
To automatically predict the IIS, we initially assess the zero-shot transfer capabilities of pre-trained NR-IQA models. The subpar performance of these models on IISA underscores the distinct nature of the two tasks and the need for IISA-specific training. To this end, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a straightforward yet effective approach for generating multiple weakly labeled image-IIS pairs from a single ground-truth one. WIISA leverages a unique property of the intrinsic scale that allows us the extrapolate the ground-truth IIS of an image to its downscaled versions via a piece-wise function. Our method can be easily integrated into the training of any model, as it is independent of specific training strategies or model characteristics. The experiments demonstrate that the proposed approach consistently boosts the performance of several methods from the NR-IQA domain retrained for IISA on the proposed dataset.
@@
  • We propose WIISA, an approach to generate multiple weakly labeled image-IIS pairs from a single ground-truth pair.
@@ 2 Related Work
The image intrinsic scale is fundamentally related to a physical property of camera and lens systems: their resolving power or ability to discern details in an image. This property is theoretically upper bounded by the camera’s sampling resolution. However, in practice, the resolving power is often lower due to imperfections in the lenses and camera system. Traditionally, these properties have been studied for the camera as a whole using metrics such as Perceptual Megapixels (P-MP) [7] and the closely related Modulation Transfer Function (MTF) [3], leading to the concept of “intrinsic camera resolution” [4]. However, because previous approaches characterize cameras under ideal laboratory conditions, they do not generalize well to real-world photos, which are affected by a multitude of unaccounted degradations. Therefore, for the first time, we propose to study the resolving power conditioned on individual images by subjectively measuring the IIS. This approach allows our method to generalize effectively to real-world photographs.
@@
Another line of research has focused on the effects of upscaling on image quality. Works such as [28, 37] aim to discriminate between authentic and artificially upscaled images. While valuable for detecting upscaling artifacts, such a binary classification does not offer a generalizable image quality metric. Kansy et al. [18] introduce the concept of ‘effective resolution”, i.e., the smallest size an image can be downscaled to, such that when it is upscaled back to its original dimensions the loss of detail is not perceptible. This is distinct from the IIS. Consider the example of a noisy image: the scale of its effective resolution should be high enough to preserve the noise pattern such that upscaling will not lead to noticeable detail loss relative to the original. In contrast, the IIS of the image should be low enough to make the noise not visible anymore, improving the perceived quality of the image.
@@ 3 Image Intrinsic Scale Assessment
As illustrated in the bottom section of Fig. 1 and supported by previous studies [33, 13, 8, 24, 11], the relationship between the perceived quality of an image and its presentation scale – defined as the scale at which it is displayed for subjective annotation – is generally assumed to follow a concave-down or monotonic function for scales smaller than 1 (i.e., the original size). When an image is downscaled, quality initially improves as degradation like blur and noise decrease. However, downsizing past a certain point reduces perceived quality primarily due to loss of detail and interpolation artifacts like aliasing.
To quantify the relationship between quality and scale, we introduce the IIS, i.e., the largest scale at which an image shows its highest perceived quality. The IIS stems from subjective evaluations of the quality of an image’s downscaled versions. Therefore, analogous to the MOS in IQA, the ground-truth IIS labels are determined by computing a scalar aggregate of the individual subjective judgments provided by multiple annotators. Based on the concept of IIS, we propose the IISA task, which has two main aspects:
@@
Even though the annotation study setup involves presenting a single image at a time, the IIS is determined by comparing the quality of rescaled versions of that image. Thus, the subjective side of IISA has both a no-reference and partial-reference nature (across scales). Nonetheless, the predictive side of IISA is a no-reference task, as it depends on a single input image for automatic IIS prediction.
@@
When downscaling images, different interpolation methods introduce artifacts like aliasing, blurring, or ringing, which affect the IIS. To achieve the highest quality, we use Lanczos interpolation due to its ability to preserve fine details and minimize aliasing through sinc-based filtering, resulting in higher image quality compared to bilinear or bicubic methods. Consequently, we employ this method for both the subjective and predictive aspects of IISA.
@@ 2. A2:
Figure 2: Example of the influence of the scaling factor $s$ on the quality $Q(I^{s})$ of downscaled images $I^{s}$ and the intrinsic scale $\overline{\Omega}(I^{s})$, based on Eq. 2. The quality is not evaluated below the lower bound $s_{lb}$. $\Omega(I)$ is the scale where the quality function is at its maximum, i.e., the IIS of the original image $I$ ($s=1$).
@@
IISA differs from NR-IQA in several key aspects. First, IISA offers a concrete way to maximize an image’s perceived quality by simply downscaling it to its intrinsic scale, while NR-IQA only provides information on an image’s quality at a fixed scale. Second, compared to NR-IQA, IISA is affected less by global image degradation, such as overall contrast and color defects. The downscaling operation mainly affects spatially limited defects such as blurs, compression artifacts, and noise. Third, NR-IQA lacks a definitive model for how image quality changes with scale. While studies like [13] explore the relationship between perceived quality and viewing distance, accurately understanding this change requires annotating an image (e.g., MOS) at multiple scales. The KonX dataset [33] provides ratings for three scales, but fine-grained sampling across all scales is impractical. On the contrary, for IISA, we can easily extrapolate the ground-truth IIS of an image to its downscaled versions by using Eq. 2. Fourth, quality ratings use a perceptually linear scale [6], whereas rescaling factors – underlying the IIS – are inherently non-linear. For example, despite the same absolute difference, changing the scale from 0.2 to 0.1 halves the image size, whereas decreasing it from 0.6 to 0.5 results in only a 17% drop. Fifth, the quality MOS in NR-IQA struggle with limited sensitivity, particularly when evaluating subtle quality differences in high-quality images resulting from restoration, enhancement, or compression methods. This has led the JPEG standards committee to explore methodologies capable of more granular quality assessment [31]. Our study demonstrates that the IISA significantly outperforms traditional NR-IQA quality ratings in detecting subtle differences, exhibiting several times greater sensitivity. See the supplementary for more details.
Applications:
IISA addresses applications where understanding scale-dependent quality is critical: 1) it optimizes image display and storage by identifying an image’s peak quality across scales, crucial for applications like the web, print, and gaming. This enhances user experience and saves storage or printing space by using the display media’s pixel density (e.g., PPI) to determine optimal display sizes. 2) IISA improves quality measurement sensitivity over traditional NR-IQA methods through implicit cross-scale comparative judgments, making it effective for detecting subtle quality degradations from restoration methods like denoising and super-resolution. See the supplementary material for more details on sensitivity. 3) IISA supports the creation of better low-level vision datasets. Unlike current methods that reduce diversity by discarding low-quality images, by downscaling high-resolution images to their IIS, IISA can help maintain both quality and diversity, essential for diverse and high-quality training datasets [12, 22]. See the supplementary for more details on the applications of IISA.
@@ 4.2 Dataset Analysis
Viewing conditions – including display characteristics (contrast, brightness), environmental lighting, viewing distance, and viewer acuity – significantly influence perceived image quality. Laboratory experiments tightly control these factors but consequently measure atypical conditions with limited variability, restricting the applicability of the findings. In contrast, crowdsourcing involves diverse and less controlled conditions, potentially reducing individual participant reliability. Nevertheless, when reliable group opinions are obtained through well-controlled methods, the resulting assessments generalize better. Our annotation framework has demonstrated high reliability, ensuring that despite individual variation, group aggregates accurately measure the perceptual IIS. Thus, we provide highly reliable data that better reflects real-world viewing conditions, enhancing the validity and applicability of our findings.
@@ 5.1 Pre-trained NR-IQA Models
The IIS stems from the quality ratings across all the downscaled versions of an image and mostly depends on low-level degradation – such as blur or noise – rather than on aesthetics or semantics. These reasons make IISA and NR-IQA closely related, despite their several key differences (see Sec. 3). Moreover, NR-IQA methods assess image quality by evaluating both the distortions and the high-quality details present in images, making them suitable for IISA. Therefore, we evaluate the performance of pre-trained NR-IQA models for zero-shot IIS prediction.
We consider TOPIQ [5], a state-of-the-art NR-IQA approach with strong generalization capabilities. We pre-train it on the SPAQ [9] and UHD-IQA [15] NR-IQA datasets, which feature high-resolution images, similar to IISA-DB. We avoid using KonIQ-10k [14] due to image overlaps with our dataset. We then apply the pre-trained TOPIQ models to predict quality scores for IISA-DB, treating such scores as IIS estimates since both metrics range in $[0,1]$. We follow the evaluation protocol described in Sec. 6.1 and present the results in Tab. 1. Both models achieve unsatisfactory performance, highlighting the inefficacy of pre-trained NR-IQA models for zero-shot IIS prediction. This emphasizes the need for IISA-specific training and techniques.
@@ 5.2 Proposed Approach: WIISA
Formally, let $I_{1}$ and $\Omega(I_{1})$ be an image and its corresponding ground-truth IIS, respectively. We randomly sample $n_{wl}$ scales in the range $[\max(\Omega(I_{1}),\delta),1]$, where $\delta$ is a threshold hyperparameter empirically set to 0.65. In our experiments, we use ${n}_{wl}=2$ unless stated otherwise. Since the scales are larger than $\Omega(I_{1})$, they fall in the second branch of Eq. 2. Then, we downscale $I_{1}$ based on each of these scales using Lanczos interpolation. Therefore, given $N=1+n_{wl}$ and the set of randomly sampled scales $\mathcal{S}=\{s_{i}\mid i=2,\dots,N\}$, we get a corresponding set of downscaled versions of the same image =$\mathcal{I}=\{I_{i}\mid i=2,\dots,N\}$. Finally, we leverage Eq. 2 to obtain the corresponding set of IIS $\overline{\mathcal{S}_{\Omega}}=\{\overline{\Omega}(I_{i})\mid i=2,\dots,N\}$. Thus, given $i=2,\dots,N$, $I_{i}\in\mathcal{I}$ represents the image obtained by rescaling $I_{1}$ to a scale $s_{i}\in\mathcal{S}$ (i.e., $I_{1}^{s_{i}}$), with an associated IIS of $\overline{\Omega}(I_{i})\in\overline{\mathcal{S}_{\Omega}}$. In the end, starting from a single ground truth image-IIS pair ($I_{1}$, $\Omega(I_{1})$), thanks to our approach, we get $n_{wl}$ weakly labeled pairs ($I_{i}$, $\overline{\Omega}(I_{i})$), with $i=2,\dots,N$. Fig. 4 shows an overview of the proposed strategy. During training, we use WIISA to augment each batch of size $B$ with additional $B\cdot n_{wl}$ weakly labeled training samples. Since our approach only requires image downscaling to generate each weak label, it can be easily applied online during training.

### v3

@@ Abstract
Image Quality Assessment (IQA) measures and predicts perceived image quality by human observers. Although recent studies have highlighted the critical influence that variations in the scale of an image have on its perceived quality, this relationship has not been systematically quantified. To bridge this gap, we introduce the Image Intrinsic Scale (IIS), defined as the largest scale where an image exhibits its highest perceived quality. We also present the Image Intrinsic Scale Assessment (IISA) task, which involves subjectively measuring and predicting the IIS based on human judgments. We develop a subjective annotation methodology and create the IISA-DB dataset, comprising 785 image-IIS pairs annotated by experts in a rigorously controlled crowdsourcing study with verified reliability. Furthermore, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a strategy that leverages how the IIS of an image varies with downscaling to generate weak labels. Experiments show that applying WIISA during the training of several IQA methods adapted for IISA consistently improves the performance compared to using only ground-truth labels. The code, dataset, and pre-trained models are available at https://github.com/SonyResearch/IISA.
[…]
@@ 1 Introduction
Most IQA datasets annotate quality ratings – in the form of Mean Opinion Scores (MOS) – for images at a fixed resolution [14, 15, 9]. To the best of our knowledge, KonX [34] was the first dataset that provided quality annotations for each image rescaled across multiple resolutions. The authors showed that perceived quality varies with image resolution (i.e., size in pixels). When an image is downscaled, its resolution decreases. Since images are typically displayed at a consistent pixel density on any given display medium, downscaled images also appear physically smaller. The perceived quality of smaller images often increases because degradation becomes less noticeable, e.g., noise grain becomes smaller or fades. Conversely, relevant high-frequency details may get lost when the resolution is too low, leading to reduced perceived quality. The former phenomenon is exemplified in Fig. 1 (top): an image that appears blurry at its original size ($scale=1$) becomes crisp and clear when downscaled to $scale=0.25$, such as the bird’s feathers. However, at a smaller scale ($scale=0.1$), fine details such as individual barbs are lost.
@@
To automatically predict the IIS, we initially assess the zero-shot transfer capabilities of pre-trained NR-IQA models. The subpar performance of these models on IISA underscores the distinct nature of the two tasks and the need for IISA-specific training. To this end, we propose WIISA (Weak-labeling for Image Intrinsic Scale Assessment), a straightforward yet effective approach for generating multiple weakly labeled image-IIS pairs from a single ground-truth one. WIISA leverages a unique property of the intrinsic scale that allows us to extrapolate the ground-truth IIS of an image to its downscaled versions via a piece-wise function. Our method can be easily integrated into the training of any model, as it is independent of specific training strategies or model characteristics. The experiments demonstrate that the proposed approach consistently boosts the performance of several methods from the NR-IQA domain retrained for IISA on the proposed dataset.
@@
  • We propose WIISA to generate multiple weakly labeled image-IIS pairs from a single ground-truth pair.
@@ 2 Related Work
The intrinsic scale is fundamentally related to a physical property of camera and lens systems: their resolving power or ability to discern details in an image. This property is theoretically limited by the camera’s sampling resolution, but in practice, it is often lower due to imperfections in the lenses and imaging system. Traditionally, these properties have been studied for the camera as a whole using metrics such as Perceptual Megapixels (P-MP) [7] and the closely related Modulation Transfer Function (MTF) [3], leading to the concept of “intrinsic camera resolution” [4]. However, because previous approaches characterize cameras under ideal laboratory conditions, they do not generalize well to real-world photos, which are affected by a multitude of unaccounted degradations. Therefore, for the first time, we propose to study the resolving power conditioned on individual images by subjectively measuring the IIS. This approach allows our method to generalize effectively to real-world photographs.
@@
Another line of research has focused on the effects of upscaling on image quality. Works such as [29, 38] aim to discriminate between authentic and artificially upscaled images. While valuable for detecting upscaling artifacts, such a binary classification does not offer a generalizable image quality metric. Kansy et al. [19] introduce the concept of “effective resolution”, i.e., the smallest size an image can be downscaled to, such that when it is upscaled back to its original dimensions the loss of detail is not perceptible. This is distinct from the IIS. Consider the example of a noisy image: the scale of its effective resolution should be high enough to preserve the noise pattern such that upscaling will not lead to noticeable detail loss relative to the original. In contrast, the IIS of the image should be low enough to make the noise not visible anymore, improving the perceived quality of the image.
@@ 3 Image Intrinsic Scale Assessment
As illustrated in Fig. 1 (bottom) and supported by previous studies [34, 13, 8, 25, 11], the relationship between the perceived quality of an image and its presentation scale is generally assumed to follow a concave-down or monotonic function for scales smaller than 1 (i.e., the original size). The “presentation scale” is the scale at which the image is displayed for subjective annotation. Downscaling an image initially improves quality by reducing degradations like blur and noise. However, excessive downscaling diminishes quality due to detail loss and artifacts like aliasing.
To quantify the relationship between quality and scale, we introduce the IIS, i.e., the largest scale at which an image shows its highest perceived quality. The IIS stems from subjective evaluations of the quality of an image’s downscaled versions. Therefore, analogous to the MOS in IQA, ground-truth IIS labels are determined by aggregating the individual subjective judgments from multiple annotators into a single scalar value. Based on the concept of IIS, we propose the IISA task, which has two main aspects:
@@
Although our annotation study presents a single image at a time, the IIS is determined by comparing the quality of rescaled versions of an image. Thus, the subjective side of IISA has a mixed nature, including comparisons across scales. In contrast, the predictive task is strictly no-reference, using a single image to predict the IIS.
@@
When downscaling images, different interpolation methods introduce artifacts like aliasing, blurring, or ringing, which affect the IIS. To achieve the highest quality, we use Lanczos interpolation. Its ability to preserve fine details and minimize aliasing via sinc-based filtering results in higher image quality compared to bilinear or bicubic methods. Consequently, we employ this method for both the subjective and predictive aspects of IISA.
@@ 2. A2:
Figure 2: Example of the influence of the scaling factor $s$ on the quality $Q(I^{s})$ of downscaled images $I^{s}$ and the intrinsic scale $\overline{\Omega}(I^{s})$, based on Eq. 2. The quality is not evaluated below the lower bound $s_{lb}$. $\Omega(I)$ is the scale where the quality function is at its maximum, i.e., the IIS of the original image II ($s=1$).
@@
IISA differs from NR-IQA in several key aspects. First, IISA offers a concrete way to maximize an image’s perceived quality by simply downscaling it to its intrinsic scale. In contrast, NR-IQA only provides information on an image’s quality at a fixed scale. Second, compared to NR-IQA, IISA is affected less by global image degradation, such as overall contrast and color defects. The downscaling operation mainly affects spatially limited defects such as blurs, compression artifacts, and noise. Third, NR-IQA lacks a definitive model for how image quality changes with scale. While studies like [13] explore the relationship between perceived quality and viewing distance, accurately understanding this change requires annotating an image (e.g., MOS) at multiple scales. The KonX dataset [34] provides ratings for three scales, but fine-grained sampling across all scales is impractical. On the contrary, for IISA, we can easily extrapolate the ground-truth IIS of an image to its downscaled versions by using Eq. 2. Fourth, quality ratings use a perceptually linear scale [6], whereas rescaling factors underlying the IIS are inherently non-linear. For example, despite the same absolute difference, changing the scale from 0.2 to 0.1 halves the image size, whereas decreasing it from 0.6 to 0.5 results in only a 17% drop. Fifth, the quality MOS in NR-IQA struggle with limited sensitivity. This is particularly true when evaluating subtle quality differences in high-quality images resulting from restoration, enhancement, or compression methods. This has led the JPEG standards committee to explore techniques capable of more granular quality assessment [32]. In the supplementary material, we prove that IISA significantly outperforms traditional NR-IQA quality ratings in detecting subtle differences, exhibiting several times greater sensitivity.
Applications
IISA addresses applications where understanding scale-dependent quality is critical: 1) it optimizes image display and storage by identifying an image’s peak quality across scales, crucial for applications like the web, print, and gaming. This enhances user experience and saves storage or printing space by using the display media’s pixel density (e.g., PPI) to determine optimal display sizes. 2) IISA improves quality measurement sensitivity over traditional NR-IQA methods because it relies on implicit cross-scale comparative judgments. This makes it effective for detecting subtle quality degradations from restoration methods like denoising and super-resolution. See the supplementary material for more details on sensitivity. 3) IISA supports the creation of better low-level vision datasets. Unlike current methods that reduce diversity by discarding low-quality images, by downscaling high-resolution images to their IIS, IISA can help maintain both quality and diversity, essential for diverse and high-quality training datasets [12, 23]. See the supplementary material for more details on the applications.
@@ 4.2 Dataset Analysis
Viewing conditions – including display characteristics (contrast, brightness), environmental lighting, viewing distance, and viewer acuity – significantly influence perceived image quality. Laboratory experiments tightly control these factors but consequently measure atypical conditions with limited variability, restricting the applicability of the findings. In contrast, crowdsourcing involves diverse and less controlled conditions, potentially reducing individual participant reliability. Nevertheless, when reliable group opinions are obtained through well-controlled methods, the resulting assessments generalize better. Our annotation framework has demonstrated high reliability, ensuring that, despite individual variation, group aggregates accurately measure the perceptual IIS. Thus, we provide highly reliable data that better reflects real-world viewing conditions, enhancing the validity and applicability of our findings.
@@ 5.1 Pre-trained NR-IQA Models
The IIS is derived from quality ratings across all downscaled versions of an image. It primarily depends on low-level degradations like blur and noise, not aesthetics or semantics. These reasons make IISA and NR-IQA closely related, despite their several key differences (see Sec. 3). Moreover, since NR-IQA methods evaluate both distortions and high-quality details, they are plausible candidates for the IISA task. Therefore, we evaluate the performance of pre-trained NR-IQA models for zero-shot IIS prediction.
We consider TOPIQ [5], a state-of-the-art NR-IQA approach with strong generalization capabilities. We pre-train it on the SPAQ [9] and UHD-IQA [15] NR-IQA datasets, which feature high-resolution images, similar to IISA-DB. We avoid using KonIQ-10k [14] due to image overlaps with our dataset. We apply the pre-trained TOPIQ models to predict quality scores on IISA-DB, which we directly use as IIS estimates as both metrics range in $[0,1]$. We follow the evaluation protocol described in Sec. 6.1 and present the results in Tab. 1. Both models achieve unsatisfactory performance, highlighting the inefficacy of pre-trained NR-IQA models for zero-shot IIS prediction. This emphasizes the need for IISA-specific training and techniques.
@@ 5.2 Proposed Approach: WIISA
Formally, let $I_{1}$ and $\Omega(I_{1})$ be an image and its corresponding ground-truth IIS, respectively. We randomly sample $n_{wl}$ scales in the range $\max(\Omega(I_{1}),\delta),1]$, where $\delta$ is a threshold hyperparameter empirically set to 0.65. In our experiments, we use ${n}_{wl}=2$ unless stated otherwise. Since the scales are larger than $\Omega(I_{1})$, they fall in the second branch of Eq. 2. Then, we downscale $I_{1}$ based on each of these scales using Lanczos interpolation. Therefore, given $N=1+n_{wl}$ and the set of randomly sampled scales $\mathcal{S}=\{s_{i}\mid i=2,\dots,N\}$, we get a corresponding set of downscaled versions of the same image $\mathcal{I}=\{I_{i}\mid i=2,\dots,N\}$. Finally, we leverage Eq. 2 to obtain the corresponding set of IIS $\overline{\mathcal{S}_{\Omega}}=\{\overline{\Omega}(I_{i})\mid i=2,\dots,N\}$. Thus, given $i=2,\dots,N$, $I_{i}\in\mathcal{I}$ represents the image obtained by rescaling $I_{1}$ to a scale $s_{i}\in\mathcal{S}$ (i.e., $I_{1}^{s_{i}}$), with an associated IIS of $\overline{\Omega}(I_{i})\in\overline{\mathcal{S}_{\Omega}}$. In the end, starting from a single ground truth image-IIS pair ($I_{1}$, $\Omega(I_{1})$), thanks to our approach, we get $n_{wl}$ weakly labeled pairs ($I_{i}$, $\overline{\Omega}(I_{i})$), with $i=2,\dots,N$. Fig. 4 shows an overview of the proposed strategy. During training, we use WIISA to augment each batch of size $B$ with additional $B\cdot n_{wl}$ weakly labeled training samples. Since our approach only requires image downscaling to generate each weak label, it can be easily applied online during training.